solve 2x^2-11x-6=0
Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(75887)

solve 2x^2-11x-6=0
2x^2-12x+x-6 = 0
2x(x-6)+(x-6)=0
(x-6)(2x+1)=0
x=6 or x=-1/2
=================
Cheers,
Stan H.

Answer by Fombitz(32388)